A package for using the bbm fonts in math environment

Torsten Hilbrich^{∗}

Printed March 6, 1996

1 Introduction

Did you ever write mathematical text and needed a character specifying the set of natural numbers? One opportunity is to use the bbold font of AMS. But this is rather an outlined than a double-striked font.

I found some fonts, called bbm which are available in roman, sans serif and typewrite type and look like those you would write on paper, double-striked left side and normal right side.

2 How to use these fonts?

You simple have to input the package bbm by typing the following:

 \usepackage{bbm}

\mathbbm The fonts can now be used in math environment by typing $\mathbbm{N}$ for getting the symbol for natural numbers: ℕ. This is the same methode like for getting a calligraphic N where you use $\mathcal{N}$.

The characters can be uses as index or superscript too. Let’s see: M_{𝕚} was created with the following sequence.

 $M_\mathbbm{i}$

Do you prefer a sans serif font for sets, or even a typewrite style? No problem,

\mathbbmss the commands \mathbbmss and \mathbbmtt do the same like \mathbbm except of

\mathbbmtt using the specified font.

Examples:

 \mathbbm{N}  ℕ
 \mathbbmss{N} ℕ
 \mathbbmtt{N} �

∗hilbrich@informatik.hu-berlin.de

2.1 What about bold symbols?

By typing \mathbold or \mathversion{bold} you switch to the bold variant of some math symbols. The selection of math version must be done outside the math environment. 2 of the fonts described above are available in bold extended series too, the roman and sans serif family. If you specify \mathbold before using the symbols, you can use bold letters. Let’s see the examples again, now in bold

version: \mathbbm{N}  � \mathbbmss{N} � \mathbbmtt{N} �

3 Where to get the fonts?

The fonts can be found at CTAN^{1}, the directory is /tex-archive/fonts/cm/bbm. You need the whole directory the fonts are located in.

1e.g. ftp.dante.de